the average age of five members family is 18 years, if the age of grandfather is included the average increases by 12 years. what is the age of grandfathers?
Answer by Cromlix(4381)

Average age of 5 members is 18 years
Therefore total ages = 5 * 18 = 90
With grandfather average increases by 12 years
= 30
So 30 * 6 = 180
180 - 90 = 90
So Grandfather is 90 years old.
Hope this helps.
:-)